# Learning Minimal NAP Specifications for Neural Network Verification

### v2

@@ Appendix A Proofs of simple approaches
Theorem A.2 (Property of Coarsen).
@@
Appendix B Proofs of statistical approaches
@@
Theorem B.1 (Property of Sample_Refine).
@@
Considering the size of the minimal specification $s$, Sample_Refine will execute $s$ iterations. If we sample $k$ times in each iteration, then the probability of selecting a mandatory neuron is at least $1-\frac{\delta}{s}$. Consequently, by applying a union bound, the algorithm will identify a NAP specification with a probability of at least $1-\delta$.
Now, let’s delve deeper into one iteration of the process. The fundamental concept is that a mandatory neuron $m$ demonstrates a stronger correlation with proving the robustness query ($\mathcal{V}(P)=1$) compared to a non-mandatory one. This enhanced correlation increases the probability of its selection significantly when $k$ is sufficiently large.
Let’s revisit the notion of $M$, representing the set of mandatory neurons with a size of $|M|$. Now, let’s focus on a specific mandatory neuron $n^{+}\in M$. We define $B_{j-}$ as the event indicating $k_{n^{-}}>k_{n^{+}}$, and $B$ as the event where $B_{j-}$ holds for any non-mandatory neuron $n^{-}\in N\setminus M$. Importantly, if $B$ fails to occur, then the algorithm will correctly identify a mandatory neuron. Hence, our primary objective is to establish that $Pr(B)\leq\frac{\delta}{s}$. Initially, employing a union bound provides:
@@
Our attention now shifts towards deriving a lower bound for $\epsilon$, which intuitively reflects the discrepancy (in terms of correlation with proving the robustness query) between a mandatory neuron and a non-mandatory one. It’s noteworthy that $Pr(P_{n}\in{\{\mathbf{1},\mathbf{0}}\})=\theta$ for any $n\in N$. Furthermore, since $n^{-}$ is non-mandatory, we can assert that $Pr(\mathcal{V}(P)=1|P_{n^{-}}\in\{\mathbf{1},\mathbf{0}\})=Pr(\mathcal{V}(P)=1)$. By leveraging these observations, we can express:
@@
Let’s view $C$ as the collection of minimal NAP specifications. We can treat $C$ as a set of clauses in a Disjunctive Normal Form (DNF) formula: $\mathcal{V}(P;C)=\neg\bigvee_{c\in C}\bigwedge_{n\in c}P_{n}$, where we explicitly specify the dependence of $\mathcal{V}$ on the clauses $C$. For example, if $C={{1,2},{3}}$, it corresponds to $\mathcal{V}(P)=\neg[(P_{1}\land P_{2})\lor P_{3}]$. Now, let $C_{j}={c\in C:n\in c}$ denote the clauses containing $n$. We reformulate $Pr(\mathcal{V}(P)=1)$ as the sum of two components: one originating from the mandatory neuron $n^{+}$ and the other from the non-mandatory neuron:
@@
Calculating $Pr(\mathcal{V}(P)=1|P_{n^{+}}\in\{\mathbf{1},\mathbf{0}\})$ follows a similar process. The only distinction arises from conditioning on $P_{n^{+}}\in\{\mathbf{1},\mathbf{0}\}$, which introduces an extra factor of $\frac{1}{\theta}$ in the first term because conditioning divides by $Pr(P_{n^{+}})=\theta$. The second term remains unchanged since no $c\notin C_{n^{+}}$ is mandatory in $P_{n^{+}}$. Substituting these two outcomes back into the equation yields:
@@
Now, our objective is to establish a lower bound for (22) across all possible $\mathcal{V}$ (equivalently, $C$), where $n^{+}$ is permitted to be mandatory in $C$. Interestingly, the worst possible $C$ can be obtained by either having $|M|$ disjoint clauses ($C={{n}:n\in M}$) or a single clause ($C={M}$ if $s=|M|$). The intuition behind this is that when $C$ consists of $|M|$ clauses, there are numerous possibilities ($|M|-1$ of them) for some $c\notin C_{n^{+}}$ to be true, making it challenging to determine that $n^{+}$ is a mandatory neuron; in such cases, $\epsilon=(1-\theta)\theta(1-\theta)^{|M|-1}$. Conversely, if $C$ comprises a single clause, then letting this clause be true becomes exceedingly challenging; in this scenario, $\epsilon=(1-\theta)\theta^{|M|}$.
@@
Theorem B.2 (Property of Sample_Coarsen).
@@
where $P$ is the sampled NAP. By assumption, there exists a NAP $P^{S}\preccurlyeq P$ of size $s$ that passes the verification. Define $G(P)=\neg(P^{S}\preccurlyeq P)$, which is 0 when $P^{S}$ subsumes $P$, i.e., all mandatory neurons in $P^{S}$ shows up in the sampled $P$. This follows that $Pr(G(P)=0)=Pr(P^{S}\preccurlyeq P)=\theta^{s}$. Note that $G(P)\geq\mathcal{V}(P)$, as the NAP $P^{S}$ suffices to prove the robustness query. We can estimate the upper bound of $\mathcal{C}(P^{L})$ by replacing $\mathcal{V}$ with $G$:

### v4

@@ Appendix D Proofs of simple approaches
Theorem D.2 (Property of Coarsen).
@@
Appendix E Proofs of statistical approaches
@@
Theorem E.1 (Property of Sample_Refine).
@@
Considering the size of the minimal specification $s$, Sample_Refine will execute $s$ iterations. If we sample $k$ times in each iteration, then the probability of selecting a essential neuron is at least $1-\frac{\delta}{s}$. Consequently, by applying a union bound, the algorithm will identify a NAP specification with a probability of at least $1-\delta$.
Now, let’s delve deeper into one iteration of the process. The fundamental concept is that a essential neuron $m$ demonstrates a stronger correlation with proving the robustness query ($\mathcal{V}(P)=1$) compared to a non-essential one. This enhanced correlation increases the probability of its selection significantly when $k$ is sufficiently large.
Let’s revisit the notion of $M$, representing the set of essential neurons with a size of $|M|$. Now, let’s focus on a specific essential neuron $n^{+}\in M$. We define $B_{j-}$ as the event indicating $k_{n^{-}}>k_{n^{+}}$, and $B$ as the event where $B_{j-}$ holds for any non-essential neuron $n^{-}\in N\setminus M$. Importantly, if $B$ fails to occur, then the algorithm will correctly identify a essential neuron. Hence, our primary objective is to establish that $Pr(B)\leq\frac{\delta}{s}$. Initially, employing a union bound provides:
@@
Our attention now shifts towards deriving a lower bound for $\epsilon$, which intuitively reflects the discrepancy (in terms of correlation with proving the robustness query) between a essential neuron and a non-essential one. It’s noteworthy that $Pr(P_{n}\in{\{\mathbf{1},\mathbf{0}}\})=\theta$ for any $n\in N$. Furthermore, since $n^{-}$ is non-essential, we can assert that $Pr(\mathcal{V}(P)=1|P_{n^{-}}\in\{\mathbf{1},\mathbf{0}\})=Pr(\mathcal{V}(P)=1)$. By leveraging these observations, we can express:
@@
Let’s view $C$ as the collection of minimal NAP specifications. We can treat $C$ as a set of clauses in a Disjunctive Normal Form (DNF) formula: $\mathcal{V}(P;C)=\neg\bigvee_{c\in C}\bigwedge_{n\in c}P_{n}$, where we explicitly specify the dependence of $\mathcal{V}$ on the clauses $C$. For example, if $C={{1,2},{3}}$, it corresponds to $\mathcal{V}(P)=\neg[(P_{1}\land P_{2})\lor P_{3}]$. Now, let $C_{j}={c\in C:n\in c}$ denote the clauses containing $n$. We reformulate $Pr(\mathcal{V}(P)=1)$ as the sum of two components: one originating from the essential neuron $n^{+}$ and the other from the non-essential neuron:
@@
Calculating $Pr(\mathcal{V}(P)=1|P_{n^{+}}\in\{\mathbf{1},\mathbf{0}\})$ follows a similar process. The only distinction arises from conditioning on $P_{n^{+}}\in\{\mathbf{1},\mathbf{0}\}$, which introduces an extra factor of $\frac{1}{\theta}$ in the first term because conditioning divides by $Pr(P_{n^{+}})=\theta$. The second term remains unchanged since no $c\notin C_{n^{+}}$ is essential in $P_{n^{+}}$. Substituting these two outcomes back into the equation yields:
@@
Now, our objective is to establish a lower bound for (22) across all possible $\mathcal{V}$ (equivalently, $C$), where $n^{+}$ is permitted to be essential in $C$. Interestingly, the worst possible $C$ can be obtained by either having $|M|$ disjoint clauses ($C={{n}:n\in M}$) or a single clause ($C={M}$ if $s=|M|$). The intuition behind this is that when $C$ consists of $|M|$ clauses, there are numerous possibilities ($|M|-1$ of them) for some $c\notin C_{n^{+}}$ to be true, making it challenging to determine that $n^{+}$ is a essential neuron; in such cases, $\epsilon=(1-\theta)\theta(1-\theta)^{|M|-1}$. Conversely, if $C$ comprises a single clause, then letting this clause be true becomes exceedingly challenging; in this scenario, $\epsilon=(1-\theta)\theta^{|M|}$.
@@
Theorem E.2 (Property of Stochastic_Coarsen).
@@
where $P$ is the sampled NAP. By assumption, there exists a NAP $P^{S}\preccurlyeq P$ of size $s$ that passes the verification. Define $G(P)=\neg(P^{S}\preccurlyeq P)$, which is 0 when $P^{S}$ subsumes $P$, i.e., all essential neurons in $P^{S}$ shows up in the sampled $P$. This follows that $Pr(G(P)=0)=Pr(P^{S}\preccurlyeq P)=\theta^{s}$. Note that $G(P)\geq\mathcal{V}(P)$, as the NAP $P^{S}$ suffices to prove the robustness query. We can estimate the upper bound of $\mathcal{C}(P^{L})$ by replacing $\mathcal{V}$ with $G$: